OpenStudy (anonymous):

A transverse wave is described by y(x,t) = 0.0065sin(3.57x – 27.7t + π/4) (Units are standard units.) Determine: a.) Amplitude b.) Wavelength c.) Frequency d.) Period e.) Speed of propagation f.) Direction of travel g.) Phase Constant h.) The y-position of a point at x = 1.50m and t = 0.12 s. (remember radians!)

OpenStudy (anonymous):

A transverse wave is normally given in the form: \[y(x,t)=Asin(kx-\omega t +\phi)\], where - A = amplitude - k = wavenumber - \(\omega\) = angular frequency - \(\phi\) = phase shift.

OpenStudy (anonymous):

\[A = 0.0065\]\[\kappa = 3.57\]\[\omega = 27.7\]\[\phi = \pi/4\]

OpenStudy (anonymous):

the frequency is f = 2w The period is T=1/f

OpenStudy (anonymous):

wave length is 2pi/k